You own a portfolio that has $2848 invested in Stock A and $3385 invested in Stock B. If the expected returns on these stocks are 11 percent and 17 percent, respectively, what is the expected return (in percent) on the portfolio?
Answer #1

Return on Stock A = 11%*2848 =313.28

Return on Stock B = 17%*3385 = 575.45

Total return on portfolio = 888.73

Total investment =2848+3385 = 6233

Expected return on porfolio =888.73/6233 = 14.26%